please show how you got the answer
Find B. Ho: M = 7 us returnel H.M= 7.2 if Seuze $=0.567 x=0.05 n=47, X= 7.112 /
Answer #1

Type II Error is failing to reject the null when it is false

\beta is the probability of type II error.

H:H= 7

H_a: \mu \neq 7

significance level = g= 0,05

This is a two-tailed test. We will fail to reject the null (commit a Type II error) if we get a Z statistic greater than -1.96 and less than 1.96.

This -1.96 & 1.96  Z-critical value corresponds to some X critical value ( X critical), such that

\frac{X_{c_1}-7}{\frac{0.567}{\sqrt{47}}}=1.96

X_{c_1}=7.162

\frac{X_{c_2}-7}{\frac{0.567}{\sqrt{47}}}=-1.96

X_{c_2}=6.838

So I will incorrectly fail to reject the null as long as a draw a sample mean that greater than 6.838 and less than 7.162.

To complete the problem what I now need to do is compute the probability of drawing a sample mean greater than 6.838 and less than 7.162 given \mu = 7.2

Thus, the probability of a Type II error is given by

P(\frac{6.838-7.2}{\frac{0.567}{\sqrt{47}}}<Z<\frac{7.162-7.2}{\frac{0.567}{\sqrt{47}}})

P(-4.377<Z<-0.459)=P(Z<-0.459)-P(Z<-4.377)

P(-4.377<Z<-0.459)=P(Z<-0.459)-P(Z<-4.377)=0.3231-0

the probability of a Type II error is \beta=0.3231